Let P(x):x2/2=x


Find the following then identify their truth values.


P (1)


P (2)


∀n,P(n)


∃n,P (n)



1
Expert's answer
2022-03-30T13:30:01-0400

"x^2\/2=x"

P(1)=1/2=1 - false

P(2)=4/2=2 truth

"\u2200n,P(n)" false

"\u2203n,P (n)" truth